# Bessel Functions and Analysis of Circular Waveguides

### v1

@@ Abstract
AMS classification:
78A50, 35Q61
Acknowledgments
L. Demkowicz, J. Mora-Paz and S. Henneking were supported by AFOSR grant FA9550-23-1-0103.
@@ 1 Introduction
The present work started as part of an initiative to perform a computational verification of our three-dimensional circularly bent fiber finite element model’s methodology [mora2025bent], implemented in hp3D [henneking2024hp3d], based on a vectorial envelope formulation [henneking2025envelope] of a time-harmonic Maxwell fiber model [henneking2021fiber, henneking2022parallel]. To do so, we consider a simplified, circularly bent, two-dimensional (slab) waveguide that has analytic solutions (derived and computed below), even for very slight bends (very large bend radii). This coiled slab waveguide model does not include the photoelastic effect [fiber1978butter, nye1985physical, rosello2016measurement], where the bending induces a stress/strain that alters the fiber’s index of refraction, since this would inhibit the analytic results that we want to compare against. Rather, the main coiling effect on the waveguide arises from the geometry of the circular bend [marcuse1976curvature, marcuse1982influence, faustini1997bend], and this is accounted for in our bent slab waveguide problem.
@@
    $\frac{1}{r}\frac{{\partial}}{{\partial}r}\left(r\frac{{\partial}p}{{\partial}r}\right)+\frac{1}{r^{2}}\frac{{\partial}^{2}p}{{\partial}\theta^{2}}+k_{0}^{2}n^{2}p=0\,$, (1.1)
@@
    $r\frac{d}{dr}\left(r\frac{du}{dr}\right)+\left(\kappa^{2}r^{2}-\lambda\right)u=0\ ,$ (1.2)
@@
    $\widetilde{r}\frac{d}{d\widetilde{r}}\left(\widetilde{r}\frac{du}{d\widetilde{r}}\right)+\left(\widetilde{r}^{\,2}-\lambda\right)u=0\ .$ (1.3)
@@
    $\frac{{\partial}u}{{\partial}r}=0\,;$ (1.5)
@@
    $\frac{{\partial}u}{{\partial}r}+ik_{0}du=0\,.$ (1.6)
@@
REMARK 1  There are two possibilities to physically interpret problem (1.1) and its boundary conditions. On one hand, $p$ may represent the pressure in time-harmonic acoustical waves, with a sound hard BC on the inner boundary. Alternatively, a derivation from Maxwell’s equations, under the assumptions of no free charges, no material magnetic susceptibility, and a linearly polarized electric field in a direction perpendicular to the plane $(r,\theta)$, yields an identical scalar eigenproblem, where $p$ corresponds to such a component of the time-harmonic electric field, $\mathbf{E}$. Its BC on the left-hand end is interpreted as a magnetic BC. Additionally, the impedance BC in the scalar problem can be shown to be equivalent to the impedance BC for electromagnetics. In the latter, the tangential components of $\mathbf{E}$ and the magnetic field $\mathbf{H}$ hold the following relationship:
$Z_{\text{clad}}\mathbf{n}\times(\mathbf{H}\times\mathbf{n})+\mathbf{E}\times\mathbf{n}=\mathbf{0},$ (1.7)
@@ 2 Evaluation of Bessel functions of large argument and large complex order
$\frac{r}{r_{0}}\frac{d}{dr}\left(\frac{r}{r_{0}}\frac{du}{dr}\right)+\left(k_{0}^{2}n_{0}^{2}\left(\frac{r}{r_{0}}\right)^{2}-\frac{\lambda}{r_{0}^{2}}\right)u$ $=0\text{ for }r\in(r_{0}-b,r_{0}+b)\,.$ (2.1)
@@
$x\coloneqq r_{0}\ln\left(\frac{r}{r_{0}}\right),$ (2.2)
@@
$\frac{d^{2}u}{dx^{2}}+\left(\kappa^{2}\exp\left(\frac{2x}{r_{0}}\right)-\mu\right)u=0\quad\text{ for }x\in\left(r_{0}\ln\left(\frac{r-b}{r_{0}}\right),r_{0}\ln\left(\frac{r+b}{r_{0}}\right)\right)\,.$ (2.3)
@@
Computation of the rescaled Bessel functions.
@@
$(m+1)(m+2)c_{m+2}+\kappa^{2}\sum_{l=0}^{m}\frac{2^{m-l}}{r_{0}^{m-l}(m-l)!}c_{l}-\mu c_{m}=0\quad\text{ for }m\in\mathbb{N}\,.$ (2.5)
@@
  • Evaluation of particular solutions to Bessel’s equation: if $\mu$ is known, for a given $r$ we evaluate the two fundamental solutions $V_{\mu}(x(r))$ and $W_{\mu}(x(r))$, truncating the series whenever the two latest increments $|c_{m}x^{m}|$ are below the desired precision. Moreover, if $C$ and $D$ are provided, we can deliver the value of the particular solution (2.6).
@@
Derivative of the rescaled Bessel function with respect to order $\bm{\mu}$.
@@
$\frac{\partial^{2}}{\partial x^{2}}\left[\frac{\partial u}{\partial\mu}\right]+\left(\kappa^{2}\exp\left(\frac{2x}{r_{0}}\right)-\mu\right)\frac{\partial u}{\partial\mu}=u\ ,$ (2.7)
@@
$b_{0}=b_{1}=0;\qquad(m+1)(m+2)b_{m+2}+\kappa^{2}\sum_{l=0}^{m}\frac{2^{m-l}}{r_{0}^{m-l}(m-l)!}b_{l}-\mu b_{m}=c_{m}\quad\text{ for }m\in\mathbb{N}\,$, (2.8)
@@
$\frac{\partial u}{\partial\mu}=C\frac{\partial V_{\mu}}{\partial\mu}(x)+C\frac{\partial W_{\mu}}{\partial\mu}(x)\,.$ (2.9)
@@
$\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{\infty}\frac{\operatorname*{Im}(\lambda_{i})\,\operatorname*{Im}(\lambda_{j})}{|\lambda_{i}-\lambda_{j}|^{2}}<\infty\,.$ (3.1)
@@
$\alpha_{n}(\lambda_{n})\coloneqq\sqrt{\kappa^{2}-\frac{\lambda_{n}}{r_{0}^{2}}}\,$, (3.2)
@@
$G_{n}=\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}\frac{\operatorname*{Im}(\lambda_{i})\,\operatorname*{Im}(\lambda_{j})}{|\lambda_{i}-\lambda_{j}|^{2}}\,,\quad n=1,2,\ldots$ (3.3)
@@ 4 Main application: Three-layer circular waveguide eigenproblem
Next, we solve the three-layer waveguide eigenproblem in the $x$ coordinate. The solution consists of three branches corresponding to the core and the two (left and right) cladding regions. The solution in the waveguide core region will be represented as a function of $x$, with $x=0$ corresponding to $r=r_{0}$; the solution in the left cladding region, $r<r_{0}-a$, will be represented as a function of $x_{1}$, such that $x_{1}=0$ corresponds to $r=r_{0}-a$; and the solution in the right cladding region, $r>r_{0}+a$, will be represented as a function of $x_{2}$, such that $x_{2}=0$ corresponds to $r=r_{0}+a$. The logarithmic transformation of $x$ is as presented in Section 2, while for $x_{1}$ and $x_{2}$ we change the base point $r_{0}$ to $r_{0}-a$ and $r_{0}+a$, respectively:
[…]
@@
$u^{\text{left}}(0)=u^{\text{core}}(-a_{1})\qquad\text{and}\qquad u^{\text{right}}(0)=u^{\text{core}}(a_{2})\ ,$ (4.13)
@@
$\frac{du^{\text{left}}}{dx_{1}}(0)=\frac{r_{0}}{r_{0}-a}\,\frac{du^{\text{core}}}{dx}(-a_{1})\qquad\text{and}\qquad\frac{du^{\text{right}}}{dx_{2}}(0)=\frac{r_{0}}{r_{0}+a}\,\frac{du^{\text{core}}}{dx}(a_{2})\,.$ (4.14)
@@
$\frac{r_{0}-a}{r_{0}-b}\,\frac{du^{\text{left}}}{dx_{1}}(-b_{1})=0\,$, (4.15)
@@
$\frac{r_{0}+a}{r_{0}+b}\,\frac{du^{\text{right}}}{dx_{2}}(b_{2})+i\kappa_{\text{clad}}du^{\text{right}}(b_{2})=0\,.$ (4.16)
Detailed description of the eigenvalue solver for a circularly bent three-layer waveguide.
@@
$u^{\text{core}}(-a_{1})=C_{0}V^{\text{core}}_{\mu}(-a_{1})+D_{0}W^{\text{core}}_{\mu}(-a_{1})\quad\text{and}\quad\frac{du^{\text{core}}}{dx}(-a_{1})=C_{0}\frac{dV^{\text{core}}_{\mu}}{dx}(-a_{1})+D_{0}\frac{dW^{\text{core}}_{\mu}}{dx}(-a_{1}).$
@@
Parameter Symbol Value in SI Units Nondimensionalized Value
Bend radius range $r_{0}$ $0.03302\text{ m},0.3302\text{ m}] [$$\in[1300,13000]$
Core’s half-width $a$ $1.27\cdot 10^{-5}$ m $0.5$
Cladding’s half-width $b$ $1.27\cdot 10^{-4}$ m $5.0$
Core’s refractive index $n_{0}$ 1.4512 $1.4512$
Cladding’s refractive index $n_{1}$ 1.45 $1.45$
Impedance BC strength $d$ $1.45$ $1.45$
Free-space wavenumber $k_{0}$ $5.905\cdot 10^{6}$ rad/m $1.49993333460866\cdot 10^{2}$
Angular frequency $\omega$ $1.77035\cdot 10^{15}$ rad/s $1.77034921739554\cdot 10^{2}$
Vacuum permeability $\mu_{0}$ $1.25664\cdot 10^{-6}$ N/A2 $0.791582400800000$
Vacuum permittivity $\epsilon_{0}$ $8.85419\cdot 10^{-12}$ F/m $0.906838390340768$
Free-space impedance $Z$ $3.76730313412\cdot 10^{2}$ V/A $0.934293045847926$
@@
On the computational implementation and the precision digits.
@@
How to interpret the numerical results.
@@
Numerical results using the impedance BC.
@@
$\begin{split}u^{\prime\prime}+(k_{0}^{2}n^{2}-\mu_{\text{straight}})u=0,\\ [\![u(\pm a)]\!]=0,\ \ [\![u^{\prime}(\pm a)]\!]=0,\\ u^{\prime}(-b)=0,\ \ u^{\prime}(b)-idk_{0}n(b)u(b)=0.\end{split}$ (4.19)
@@
$\operatorname*{Im}$} $2.623592454862567\cdot 10^{-21}$ $4.636103134791148\cdot 10^{-28}$ $3.566233180608575\cdot 10^{-32}$
@@ 5 Replacing the impedance BC with a perfectly matched layer
    $H^{(2)}_{\beta}(k_{0}n_{\text{clad}}z)\approx\sqrt{\frac{2}{\pi k_{0}n_{\text{clad}}z}}\exp\{-i(k_{0}n_{\text{clad}}z-\beta\pi/2-\pi/4)\}$ (5.1)
@@
  • In order to implement the PML in practice, we designate the new path in the boundary layer by the formula $z=z(r)$, and appropriately modify the governing relation (1.3). It is important to realize that the projection of the solution onto power series expansions remains valid even after implementing the complex argument $z$ in the boundary layer (see Remark 5 about the additional coordinate transformation applied to $x$). The new complex-valued path $z(r)$ needs to start at a point on the real line between $r_{0}+a$ and $r_{0}+b$ and to converge smoothly to a preset point $z_{\text{PML}}$ (see Fig. 5), which has a negative imaginary part that depends on the cladding wavenumber $\kappa_{1}$ and a positive real parameter $C_{\text{PML}}$, which controls the strength of the PML absorption:
    $z_{\text{PML}}=r_{0}+b-i\,\frac{C_{\text{PML}}}{\kappa_{1}}\,$, (5.2)
@@
REMARK 2  As presented in [collino1998perfectly], Collino and Monk carried out the impact analysis on the Hankel function’s asymptotic behavior (5.1). In their case, however, they used the $H^{(1)}_{\beta}$ Hankel functions rather than $H^{(2)}_{\beta}$, as well as a PML path with positive imaginary components due to their choice of sign (convention) in the time-harmonic ansatz (which is opposite to the one used in the present work, that coincides with the ansatz in [marcuse1982light]).
REMARK 3   It is important to emphasize that $z=z_{\text{PML}}$ replaces the physical coordinate $r=r_{0}+b$, and not the transformed coordinate $x_{2}=b_{2}$ defined in (4.4). Only after transforming $r\mapsto z(r)$ in the PML region, is the transformation made to $x_{2}$:
@@
Numerical experiments using the PML BC.
@@
Table 8: Comparison of the model predicted mode propagation constants in the three-layer circularly bent waveguide for three distinct bend radii, using either the impedance BC or the PML BC, applied at $r=r_{0}+b$. The digits, or entire values, that differ between the two conditions are shaded in the impedance BC column only. The real components tend to be more aligned than the imaginary components, and, at least for this specific problem, are of the same order of magnitude. Moreover, the imaginary components are better matched between these two BCs for weaker bends (i.e. larger bend radii). Because the PML BC is a better representation of the proper radiation condition, its results are trusted to be more accurate than those found using the impedance BC.
@@
Main differences between the results with the impedance BC and the PML BC.
@@
    Again, by looking at Figure 4, we notice that the solution profile, after a few fluctuations, nearly vanishes at the right boundary. This responds to a magnitude disparity between the terms of the impedance BC (cf. (1.6), where the coefficient that multiplies $u$ is much larger than the one acting on ${\partial}u/{\partial}r$). In contrast, the mode profiles obtained with the PML BC do not present that feature (consider Fig. 6 and Fig. 7).
@@
Robustness of the predicted solution with respect to the PML parameters.

### v2

@@ Abstract
{MSCcodes}
34B30, 65L15, 78M99, 33C10
@@ 1 Introduction
The present work started as part of an initiative to perform a computational verification of our three-dimensional circularly bent fiber finite element model’s methodology [mora2025bent], implemented in hphp3D [henneking2024hp3d], based on a vectorial envelope formulation [henneking2025envelope] of a time-harmonic Maxwell fiber model [henneking2021fiber, henneking2022parallel]. To do so, we consider a simplified, circularly bent, two-dimensional (slab) waveguide that has analytic solutions (derived and computed below), even for very slight bends (very large bend radii). This coiled slab waveguide model does not include the photoelastic effect [fiber1978butter, nye1985physical, rosello2016measurement], where the bending induces a stress/strain that alters the fiber’s index of refraction, since this would inhibit the analytic results that we want to compare against. Rather, the main coiling effect on the waveguide arises from the geometry of the circular bend [marcuse1976curvature, marcuse1982influence, faustini1997bend], and this is accounted for in our bent slab waveguide problem.
@@
    (1.1) $\frac{1}{r}\frac{{\partial}}{{\partial}r}\left(r\frac{{\partial}p}{{\partial}r}\right)+\frac{1}{r^{2}}\frac{{\partial}^{2}p}{{\partial}\theta^{2}}+k_{0}^{2}n^{2}p=0\,$,
@@
    (1.2) $r\frac{d}{dr}\left(r\frac{du}{dr}\right)+\left(\kappa^{2}r^{2}-\lambda\right)u=0\ ,$
@@
    (1.3) $\widetilde{r}\frac{d}{d\widetilde{r}}\left(\widetilde{r}\frac{du}{d\widetilde{r}}\right)+\left(\widetilde{r}^{\,2}-\lambda\right)u=0\ .$
@@
    (1.5) $\frac{{\partial}u}{{\partial}r}=0\,;$
@@
    (1.6) $\frac{{\partial}u}{{\partial}r}+ik_{0}du=0\,.$
@@
Remark 1.1.
There are two possibilities to physically interpret problem (1.1) and its boundary conditions. On one hand, $p$ may represent the pressure in time-harmonic acoustical waves, with a sound hard BC on the inner boundary. Alternatively, a derivation from Maxwell’s equations, under the assumptions of no free charges, no material magnetic susceptibility, and a linearly polarized electric field in a direction perpendicular to the plane $(r,\theta)$, yields an identical scalar eigenproblem, where $p$ corresponds to such a component of the time-harmonic electric field, $\mathbf{E}$. Its BC on the left-hand end is interpreted as a magnetic BC. Additionally, the impedance BC in the scalar problem can be shown to be equivalent to the impedance BC for electromagnetics. In the latter, the tangential components of $\mathbf{E}$ and the magnetic field $\mathbf{H}$ hold the following relationship:
(1.7) $Z_{\text{clad}}\mathbf{n}\times(\mathbf{H}\times\mathbf{n})+\mathbf{E}\times\mathbf{n}=\mathbf{0},$
@@ 2 Evaluation of Bessel functions of large argument and large complex order
(2.8) $\frac{r}{r_{0}}\frac{d}{dr}\left(\frac{r}{r_{0}}\frac{du}{dr}\right)+\left(k_{0}^{2}n_{0}^{2}\left(\frac{r}{r_{0}}\right)^{2}-\frac{\lambda}{r_{0}^{2}}\right)u$ $=0\text{ for }r\in(r_{0}-b,r_{0}+b)\,.$
@@
(2.9) $x\coloneqq r_{0}\ln\left(\frac{r}{r_{0}}\right),$
@@
(2.10) $\frac{d^{2}u}{dx^{2}}+\left(\kappa^{2}\exp\left(\frac{2x}{r_{0}}\right)-\mu\right)u=0\quad\text{ for }x\in\left(r_{0}\ln\left(\frac{r-b}{r_{0}}\right),r_{0}\ln\left(\frac{r+b}{r_{0}}\right)\right)\,.$
@@
Computation of the rescaled Bessel functions
@@
(2.12) $(m+1)(m+2)c_{m+2}+\kappa^{2}\sum_{l=0}^{m}\frac{2^{m-l}}{r_{0}^{m-l}(m-l)!}c_{l}-\mu c_{m}=0\quad\text{ for }m\in\mathbb{N}\,.$
@@
[…]
@@
Derivative of the rescaled Bessel function with respect to order $\bm{\mu}$
@@
(2.14) $\frac{\partial^{2}}{\partial x^{2}}\left[\frac{\partial u}{\partial\mu}\right]+\left(\kappa^{2}\exp\left(\frac{2x}{r_{0}}\right)-\mu\right)\frac{\partial u}{\partial\mu}=u\ ,$
@@
(2.15) $b_{0}=b_{1}=0;\qquad(m+1)(m+2)b_{m+2}+\kappa^{2}\sum_{l=0}^{m}\frac{2^{m-l}}{r_{0}^{m-l}(m-l)!}b_{l}-\mu b_{m}=c_{m}\quad\text{ for }m\in\mathbb{N}\,$,
@@
(2.16) $\frac{\partial u}{\partial\mu}=C\frac{\partial V_{\mu}}{\partial\mu}(x)+D\frac{\partial W_{\mu}}{\partial\mu}(x)\,.$
@@
(3.17) $\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{\infty}\frac{\operatorname*{Im}(\lambda_{i})\,\operatorname*{Im}(\lambda_{j})}{|\lambda_{i}-\lambda_{j}|^{2}}<\infty\,.$
@@
(3.18) $\alpha_{n}(\lambda_{n})\coloneqq\sqrt{\kappa^{2}-\frac{\lambda_{n}}{r_{0}^{2}}}\,$,
@@
(3.19) $G_{n}=\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}\frac{\operatorname*{Im}(\lambda_{i})\,\operatorname*{Im}(\lambda_{j})}{|\lambda_{i}-\lambda_{j}|^{2}}\,,\quad n=1,2,\ldots$
@@ 4 Main application: Three-layer circular waveguide eigenproblem
Next, we solve the three-layer waveguide eigenproblem in the $x$ coordinate. The solution consists of three branches corresponding to the core and the two (left and right) cladding regions. The solution in the waveguide core region will be represented as a function of $x$, with $x=0$ corresponding to $r=r_{0}$; the solution in the left cladding region, $r<r_{0}-a$, will be represented as a function of $x_{1}$, such that $x_{1}=0$ corresponds to $r=r_{0}-a$; and the solution in the right cladding region, $r>r_{0}+a$, will be represented as a function of $x_{2}$, such that $x_{2}=0$ corresponds to $r=r_{0}+a$. The logarithmic transformation of $x$ is as presented in Section 2, cf. (2.9), while for $x_{1}$ and $x_{2}$ we change the base point $r_{0}$ to $r_{0}-a$ and $r_{0}+a$, respectively:
$x_{1}(r)\coloneqq(r_{0}-a)\ln\left(\frac{r}{r_{0}-a}\right),\ \ x(r)\coloneqq r_{0}\ln\left(\frac{r}{r_{0}}\right),\ \ x_{1}(r)\coloneqq(r_{0}+a)\ln\left(\frac{r}{r_{0}+a}\right).$
@@
(4.32) $u^{\text{left}}(0)=u^{\text{core}}(-a_{1})\qquad\text{and}\qquad u^{\text{right}}(0)=u^{\text{core}}(a_{2})\ ,$
@@
(4.33) $\frac{du^{\text{left}}}{dx_{1}}(0)=\frac{r_{0}}{r_{0}-a}\,\frac{du^{\text{core}}}{dx}(-a_{1})\qquad\text{and}\qquad\frac{du^{\text{right}}}{dx_{2}}(0)=\frac{r_{0}}{r_{0}+a}\,\frac{du^{\text{core}}}{dx}(a_{2})\,.$
@@
(4.34) $\frac{r_{0}-a}{r_{0}-b}\,\frac{du^{\text{left}}}{dx_{1}}(-b_{1})=0\,$,
@@
(4.35) $\frac{r_{0}+a}{r_{0}+b}\,\frac{du^{\text{right}}}{dx_{2}}(b_{2})+i\kappa_{\text{clad}}du^{\text{right}}(b_{2})=0\,.$
Detailed description of the eigenvalue solver for a circularly bent three-layer waveguide
@@
$u^{\text{core}}(-a_{1})=C_{0}V^{\text{core}}_{\mu}(-a_{1})+D_{0}W^{\text{core}}_{\mu}(-a_{1}),$
and
$\frac{du^{\text{core}}}{dx}(-a_{1})=C_{0}\frac{dV^{\text{core}}_{\mu}}{dx}(-a_{1})+D_{0}\frac{dW^{\text{core}}_{\mu}}{dx}(-a_{1}).$
@@
Parameter Symbol Value in SI Units Nondimensional Value
Bend radius range
$r_{0}$
$0.03302\text{ m},0.3302\text{ m}]$
$[1300,13000]$
Core’s half-width
$a$
$1.27\cdot 10^{-5}$ m
$0.5$
Cladding’s half-width
$b$
$1.27\cdot 10^{-4}$ m
$5.0$
Core’s refractive index
$n_{0}$
1.4512
$1.4512$
Cladding’s refractive index
$n_{1}$
1.45
$1.45$
Impedance BC strength
$d$
$1.45$
$1.45$
Free-space wavenumber
$k_{0}$
$5.905\cdot 10^{6}$ rad/m
$1.49993333460866\cdot 10^{2}$
Angular frequency
$\omega$
$1.77035\cdot 10^{15}$ rad/s
$1.77034921739554\cdot 10^{2}$
Vacuum permeability
$\mu_{0}$
$1.25664\cdot 10^{-6}$ N/A2
$0.791582400800000$
Vacuum permittivity
$\epsilon_{0}$
$8.85419\cdot 10^{-12}$ F/m
$0.906838390340768$
Free-space impedance
$Z$
$3.76730313412\cdot 10^{2}$ V/A
$0.934293045847926$
@@
On the computational implementation and the precision digits
@@
How to interpret the numerical results
@@
Remark 4.1.
As explained along the algorithm derivations above, in addition to the eigenvalue, the unspecified coefficient $D_{0}$ (or $C_{0}$ for odd modes) is an unknown that must also be solved for. Of course, in the process of computing the numerical experiments herein published, we do solve for both unknowns and get a converged solution for such a coefficient too. We choose not to report its value in the tables below, to focus more on the eigenvalue results, as the latter are of higher interest for the physical application under investigation. However, in the accompanying code, the users will get the values of $C_{0}$ and $D_{0}$ as additional outputs from the scripts that reproduce the present results.
Numerical results using the impedance BC
@@
(4.38) $\begin{split}u^{\prime\prime}+(k_{0}^{2}n^{2}-\mu_{\text{straight}})u=0,\\ [\![u(\pm a)]\!]=0,\ \ [\![u^{\prime}(\pm a)]\!]=0,\\ u^{\prime}(-b)=0,\ \ u^{\prime}(b)-idk_{0}n(b)u(b)=0.\end{split}$
@@
$\operatorname*{Im}$} $2.62359245486257\cdot 10^{-21}$ $4.63610313479115\cdot 10^{-28}$ $3.56623318060858\cdot 10^{-32}$
@@ 5 Replacing the impedance BC with a perfectly matched layer
    (5.61) $H^{(2)}_{\beta}(k_{0}n_{\text{clad}}z)\approx\sqrt{\frac{2}{\pi k_{0}n_{\text{clad}}z}}\exp\{-i(k_{0}n_{\text{clad}}z-\beta\pi/2-\pi/4)\}$
@@
  • In order to implement the PML in practice, we designate the new path in the boundary layer by the formula $z=z(r)$, and appropriately modify the governing relation (1.3). It is important to realize that the projection of the solution onto power series expansions remains valid even after implementing the complex argument $z$ in the boundary layer (see Remark 5.2 about the additional coordinate transformation applied to $x$). The new complex-valued path $z(r)$ needs to start at a point on the real line between $r_{0}+a$ and $r_{0}+b$ and to converge smoothly to a preset point $z_{\text{PML}}$ (see Fig. 5), which has a negative imaginary part that depends on the cladding wavenumber $\kappa_{1}$ and a positive real parameter $C_{\text{PML}}$, which controls the strength of the PML absorption:
    (5.62) $z_{\text{PML}}=r_{0}+b-i\,\frac{C_{\text{PML}}}{\kappa_{1}}\,$,
@@
Remark 5.1.
As presented in [collino1998perfectly], Collino and Monk carried out the impact analysis on the Hankel function’s asymptotic behavior (5.61). In their case, however, they used the $H^{(1)}_{\beta}$ Hankel functions rather than $H^{(2)}_{\beta}$, as well as a PML path with positive imaginary components due to their choice of sign (convention) in the time-harmonic ansatz (which is opposite to the one used in the present work, that coincides with the ansatz in [marcuse1982light]).
Remark 5.2.
It is important to emphasize that $z=z_{\text{PML}}$ replaces the physical coordinate $r=r_{0}+b$, and not the transformed coordinate $x_{2}=b_{2}$ defined in (4.23). Only after transforming $r\mapsto z(r)$ in the PML region, is the transformation made to $x_{2}$:
@@
Numerical experiments using the PML BC
@@
Table 8: Comparison of the model-predicted mode propagation constants in the three-layer circularly bent waveguide for three distinct bend radii, using either the impedance BC or the PML BC, applied at $r=r_{0}+b$. The digits, or entire values, that differ between the two conditions are shaded in the impedance BC column only. The real components tend to be more aligned than the imaginary components, and, at least for this specific problem, are of the same order of magnitude. Moreover, the imaginary components are better matched between these two BCs for weaker bends (i.e. larger bend radii). Because the PML BC is a better representation of the proper radiation condition, its results are trusted to be more accurate than those found using the impedance BC.
@@
Main differences between the results with the impedance BC and the PML BC
@@
    Again, by looking at Fig. 4, we notice that the solution profile, after a few fluctuations, nearly vanishes at the right boundary. This responds to a magnitude disparity between the terms of the impedance BC (cf. (1.6), where the coefficient that multiplies $u$ is much larger than the one acting on ${\partial}u/{\partial}r$). In contrast, the mode profiles obtained with the PML BC do not present that feature (consider Fig. 6 and Fig. 7).
@@
Robustness of the predicted solution with respect to the PML parameters